OpenStudy (anonymous):

3). Find the equation of the line that passes through the following pair of points.(Write your answers in slope-intercept form.) (-1, -4), (1, -2) y=?

OpenStudy (anonymous):

m = -2+4/1--1 = 2/2 = 1 y = x +b -2 = 1+b b = -3 y = x -3